# Adsorption of polyelectrolytes in the presence of varying dielectric discontinuity between solution and substrate

### v2

@@ III.3 Single PE perpendicular to the substrate
Figure 9: The normalized potential of mean force $\beta V/L$ as a function of the distance $z$ between the substrate and the center of the PE bead closest to it in $1:-1$ salt with the concentration of (a) $Z^{\rm c}c^{\rm c}=30$ mM, (b) $Z^{\rm c}c^{\rm c}=60$ mM, and (c) $Z^{\rm c}c^{\rm c}=250$ mM for varying relative dielectric constant of the substrate $\varepsilon_{\rm s}$. The vertical dashed lines represent the distance at which the PE head bead contacts the substrate surface at $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, while the dotted lines represent the distance at which they are one ion diameter apart at $z_{2}=z_{1}+\sigma^{i}$.
The electrophoretic translocation of DNA and other biomolecules through membrane nanopores is of significant importance in biotechnology and has recently attracted attention Palyulin, Ala-Nissila, and Metzler (2014); Buyukdagli, Sarabadani, and Ala-Nissila (2019). A key aspect of sequence reading during translocation is the regulation of DNA movement Clarke et al. (2009), which necessitates a quantitative characterization of DNA-medium interactions. Here, we investigate a DNA-like rod molecule oriented perpendicular to the substrate, while varying the dielectric permittivity $\varepsilon_{\rm s}$ and salt concentration. This model is relevant for studying the DNA approach stage towards the membrane when controlled by purely electrostatic forces. The PE here has a length of $10$ nm with $\lambda_{0}=-6$ e/nm. For this case, we vary the substrate dielectric constant as in the previous sections of this manuscript. The PE approach stage towards a thin membrane has been analytically studied using field-theoretical approaches beyond the MF theory to highlight the influence of dielectric mismatch on polymer translocation through membranes Buyukdagli and Ala-Nissila (2016a, b); Buyukdagli, Sarabadani, and Ala-Nissila (2018). It was shown that polymers passing through carbon-based membranes with $\varepsilon_{\rm s}=2$ face a significant Coulombic repulsion. Conversely, membranes engineered to have a high dielectric constant are strongly attractive and can trap the PE in the pore. In our system, we study the approach on the cis side of the substrate only because the substrate medium virtually fills the whole region for $z<0$.
@@ IV Summary and Conclusions
In this work, we have used a simple coarse-grained MD model to assess the spontaneous adsorption of rigid rodlike PEs on uncharged substrates in water. We demonstrated that the adsorption of PEs to the substrates can be regulated via the substrate dielectric constant, salt concentration, PE charge, and ion valency. As expected, when the substrate dielectric constant $\varepsilon_{\rm s}$ is lower than that of water $\varepsilon_{\rm w}$, PEs are repelled from the substrate due to the presence of image charges with the same sign. On the contrary, when the substrate has a high dielectric constant as compared to that of water ($\varepsilon_{\rm s}>\varepsilon_{\rm w}$), the induced image charges have opposite sign to the PEs leading to attraction between the PEs and the substrate which may facilitate adsorption.
As added salt concentration increases, the resulting behavior becomes more complex due to the interplay of steric and electrostatic forces between the substrate and the PE Drechsler et al. (2010). If the PE is strongly charged, the condensed counterion layer in the region between the PE and the substrate swells, leading to significant steric repulsion. For weakly charged PEs, the electrostatic interactions of the PE with counterions are weak, and the condensed counterion layer collapses and reduces the steric barrier between the PE and the substrate. Consequently, at the smallest distances between the substrate and the PE, the interactions are mainly governed by electrostatic and entropic depletion forces that attract the PEs to the substrate Alarcón, Pérez, and Goicochea (2013), especially at high salt concentrations.

### v3

@@ III.3 Single PE perpendicular to the substrate
The electrophoretic translocation of DNA and other biomolecules through membrane nanopores is of significant importance in biotechnology and has recently attracted attention Palyulin, Ala-Nissila, and Metzler (2014); Buyukdagli, Sarabadani, and Ala-Nissila (2019). A key aspect of sequence reading during translocation is the regulation of DNA movement Clarke et al. (2009), which necessitates a quantitative characterization of DNA-medium interactions. Here, we investigate a DNA-like rod molecule oriented perpendicular to the substrate, while varying the dielectric permittivity $\varepsilon_{\rm s}$ and salt concentration. This model is relevant for studying the DNA approach stage towards the membrane when controlled by purely electrostatic forces. The PE here has a length of $10$ nm with $\lambda_{0}=-6$ e/nm. For this case, we vary the substrate dielectric constant as in the Sec. III of this manuscript. The PE approach stage towards a thin membrane has been analytically studied using field-theoretical approaches beyond the MF theory to highlight the influence of dielectric mismatch on polymer translocation through membranes Buyukdagli and Ala-Nissila (2016a, b); Buyukdagli, Sarabadani, and Ala-Nissila (2018). It was shown that polymers passing through carbon-based membranes with $\varepsilon_{\rm s}=2$ face a significant Coulombic repulsion. Conversely, membranes engineered to have a high dielectric constant are strongly attractive and can trap the PE in the pore. In our system, we study the approach on the cis side of the substrate only because the substrate medium virtually fills the whole region for $z<0$.
Figure 9: The normalized potential of mean force $\beta V/L$ as a function of the distance $z$ between the substrate and the center of the PE bead closest to it in $1:1$ salt with the concentration of (a) $Z^{\rm c}c^{\rm c}=30$ mM, (b) $Z^{\rm c}c^{\rm c}=60$ mM, and (c) $Z^{\rm c}c^{\rm c}=250$ mM for varying relative dielectric constant of the substrate $\varepsilon_{\rm s}$. The vertical dashed lines represent the distance at which the PE head bead contacts the substrate surface at $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, while the dotted lines represent the distance at which they are one ion diameter apart at $z_{2}=z_{1}+\sigma^{i}$.
@@
Monte Carlo simulations have demonstrated that rodlike PEs in the vicinity of an oppositely-charged substrate with $\Delta\varepsilon=0$ tend to align parallel to the substrate Messina (2006). Our results support the observed behavior close to substrates with $\varepsilon_{\rm s}=8000$, where the attraction of the PE to the substrate is stronger when the PE is aligned parallel to the substrate as compared to the perpendicular orientation (cf. Figs. 2 and 9). Moreover, theoretical investigations revealed that perpendicular orientation of a PE near a like-charged substrate with $\Delta\varepsilon=0$ is more favorable than parallel orientation Buyukdagli and Podgornik (2019b, a), and this preference is further enhanced by increasing counterion valency Buyukdagli and Podgornik (2019b). Consistently, Fig. 9 shows that the repulsive force on the PE near a substrate with $\varepsilon_{\rm s}=2$ is weaker as compared to the repulsion observed for a parallel orientation (see Fig. 2). This indicates a shift in the orientation of the PE, from being perpendicular to becoming parallel to the substrate near its surface.
@@ IV Summary and Conclusions
In this work, we have used a simple coarse-grained MD model to assess the spontaneous adsorption of rigid rodlike PEs on uncharged substrates in implicit water. We demonstrated that the adsorption of PEs to the substrates can be regulated via the substrate dielectric constant, salt concentration, PE charge, and ion valency. As expected, when the substrate dielectric constant $\varepsilon_{\rm s}$ is lower than that of water $\varepsilon_{\rm w}$, PEs are repelled from the substrate due to the presence of image charges with the same sign. On the contrary, when the substrate has a high dielectric constant as compared to that of water ($\varepsilon_{\rm s}>\varepsilon_{\rm w}$), the induced image charges have opposite sign to the PEs leading to attraction between the PEs and the substrate which may facilitate adsorption.
As added salt concentration increases, the resulting behavior becomes more complex due to the interplay of steric and electrostatic forces between the substrate and the PE Drechsler et al. (2010). If the PE is strongly charged, the condensed counterion layer in the region between the PE and the substrate swells, leading to significant steric repulsion Netz and Joanny (1999a). For weakly charged PEs, the electrostatic interactions of the PE with counterions are weak, and the condensed counterion layer collapses and reduces the steric barrier between the PE and the substrate. Consequently, at the smallest distances between the substrate and the PE, the interactions are mainly governed by electrostatic and entropic depletion forces that attract the PEs to the substrate Alarcón, Pérez, and Goicochea (2013), especially at high salt concentrations.
@@
Our work has focused on the interaction of rigid PEs with planar flat substrates. In this geometry, image charge calculation results in effective surface element charges unaffected by the induced charges on neighboring elements due to the condition $\bm{n}\perp\bm{E}^{i}$. In future works, it would be interesting to consider other substrate geometries, such as convex or concave substrates interacting with flexible PEs. In such cases, the image charge calculation is influenced by the system geometry and the charge distribution Wu et al. (2018). This allows, e.g., curvature controlled creation of diverse surface charge patterns Pogharian et al. (2024). For example, in salt solutions in contact with a substrate with $\varepsilon_{\rm s}<\varepsilon_{\rm w}$, the repulsive forces on ions weaken near convex regions and strengthen near concave ones Wu et al. (2018); Solis and Olvera de la Cruz (2021). Also, a geometry change from an insulating planar substrate to a spherical shape results in a decrease of image-charge repulsion between a flexible PE and the substrate surface Cherstvy and Winkler (2012).
Our modelling results can aid in the interpretation of experimental data and enhance the performance of applications. For example, PE elastomers have widespread applications in flexible electronic devices, such as polymer-based transducers Pavlenko et al. (2016) and soft robotics Nguyen et al. (2024), but also in energy harvesting Guo and Helseth (2015), energy storage Qiu et al. (2022), and in printable electronics Tong and Tong (2022). In these applications, the dielectric mismatch between the substrate and the PE solutions affects device efficiency and performance. For example, metal oxide electrodes with higher dielectric constant significantly improve the capacitance and the electrochemical efficacy of supercapacitors Tu et al. (2020). An interesting application direction for materials with high dielectric constants is ingestible electronics and edible composites made from natural biopolymers Konwar et al. (2022). Characterization of PE-substrate interactions is also important for optimizing other applications, including biosensors Hensel et al. (2020), artificial cells Xu, Hu, and Chen (2016), drug delivery systems Zhang et al. (2018), and self-healing coatings Abu-Thabit and Hamdy (2016).